436 nm light shines on a diffraction grating with 4908 lines/cm. What is the angular separation between the second order maximum and the third order maximum?
Answer #1

Given that :

wavelength of the light, \lambda = 436 nm = 436 x 10-9 m

diffraction grating, d = 4908 lines/cm = (1 cm / 4908)

d = (1 m / 4.9 x 105)

For the second order maximum, m = 2

using an equation -

d sin \theta = m \lambda                                               

or    sin \theta = m \lambda / d

inserting the values in above eq.

sin \theta2 = (2) (436 x 10-9 m) (4.9 x 105) / (1 m)

sin \theta2 = 4272.8 x 10-4

\theta2 = sin-1 (0.42728)

\theta2 = 25.3 degree

For the third order maximum, m = 3

sin \theta3 = m \lambda / d

inserting the values in above eq.

sin \theta3 = (3) (436 x 10-9 m) (4.9 x 105) / (1 m)

sin \theta3 = 6409.2 x 10-4

\theta3 = sin-1 (0.64092)

\theta3 = 39.8 degree

angular separation between the second order maximum and the third order maximum which will be given as ::

\Delta\theta = \theta3 - \theta2

\Delta\theta = [(39.8) - (25.3)] degree

\Delta\theta = 14.5 degree